How do I solve these equations?
1/2x=1/5x+5/2
1.1x-2.7=0.7x-1.02

Answer by stanbon(75887)

1/2x=1/5x+5/2
Multiply thru by 10 to get:
5x = 2x + 25
3x = 25
x = 8 1/3
======================

1.1x-2.7=0.7x-1.02
Multiply thru by 100 to get:
110x - 270 = 70x - 102
-----
40x = 168
10x = 42
x = 4.2
====================
Cheers,
Stan H.
====================